Find the corresponding area between z = 0 and each of the following: a. z = 0.96 b. z = 2.74 c. z = -1.67


1
Expert's answer
2022-03-15T13:26:42-0400

a)

P(0<Z<0.96)=P(Z<0.96) - P(Z<0)

From the normal distribution table,

P(Z<0.96)=0.8314

P(Z<0)=0.5

Therefore,

P(0<Z<0.96) = 0.8314-0.5

=0.3314


b)

P(0<Z<2.74)=P(Z<2.74) - P(Z<0)

From the normal distribution table,

P(Z<2.74)=0.9969

P(Z<0)=0.5

P(0<Z<2.74)= 0.9969-0.5

=0.4969


c)

P(-1.67<Z<0)=P(Z<0) - P(Z<-1.67)

From the normal distribution table,

P(Z<0)= 0.5

P(Z<-1.67)= 0.0475

P(-1.67<Z<0)=0.5 - 0.0475

=0.4525